At 0°c and one atmosphere pressure a gas occupies 100cc .if the pressure is increased to one and half time and temperature is increased bye one third of absolute temperature then find the final volume of gas will be?

Grade:11

1 Answers

Shubham
6 years ago
 
Misplaced &
Þ      
P1=P
  ; 
T1=273oK
                    
P2=32P
T2=T1+T13=43×273oK
                   
V2=2P3P×43×273273×100cc=8009cc=88.888cc
                                 = 88.9 cc